# TritonRL: Training LLMs to Think and Code Triton Without Cheating

### v1

@@ 3.2 Main Experiment Results
Overall Performance on Level 1 Tasks.
The left side of Table 1 presents the performance comparison results for pass@10 evaluated with robust verifiers (syntax and functionality) on KernelBench Level 1 tasks. TritonRL consistently outperforms most baseline models with <32<32B parameter sizes in terms of validity, correctness, and speedup. In particular, TritonRL surpasses AutoTriton, which also leverages SFT and RL, by achieving higher correctness and speedup, underscoring the advantages of our hierarchical reward assignment. Notably, the correctness metric improves from 44% (SFT only) to 55% after RL, indicating that RL provides substantial gains in addition to supervised fine-tuning. Furthermore, TritonRL achieves on-par performance to much larger models, highlighting the efficiency of our approach in enabling smaller models to excel in specialized code generation tasks. We further evaluated inference scaling by varying the number of sampled attempts ($k=1,5,10$), as illustrated in Figure 4. The correctness of TritonRL increases with more samples, whereas KernelLLM and Qwen3-8B show limited improvement, suggesting that TritonRL generates a more diverse set of codes and benefits from additional sampling during inference. Additional pass@1 and pass@5 results are provided in Appendix I.2.
Figure 5: Side-by-side comparison of base models and their post-trained variants. Successful training should reduce invalid syntax errors (yellow) and functional invalidity (red), while increasing correctness (green) from individual base model. Only TritonRL shows this outcome. For each metric, the numbers represent the ratio of corresponding samples among 10 generated samples for each of the 100 tasks in KernelBench Level 1.
Effectiveness of Validity Reward.
We analyze the validity of Triton codes generated by fine-tuned models to understand the types of errors each model is prone to. To examine the effects of fine-tuning on validity, we also include the base models of TritonRL and the baseline models. In Figure 5, although both AutoTriton and TritonRL achieve relatively high rates of validity compared to KernelLLM, a more detailed breakdown reveals that AutoTriton exhibits a much higher proportion of functionally invalid codes. Interestingly, the base model of AutoTriton shows a low rate of functionally and syntactically invalid codes, indicating that the fine-tuning process of AutoTriton may have led to learn functionally invalid codes. In contrast, TritonRL generates significantly fewer invalid codes in terms of both syntax and functionality after fine-tuning, demonstrating the effectiveness of our robust verification in enhancing code quality.
Moreover, Table 1 highlights how heavily prior models relied on cheating shortcuts. Without functionality verification (w/o robust), AutoTriton’s correctness jumps from 50% to 70%, revealing its tendency to exploit reward-driven shortcuts rather than produce true Triton code. In contrast, TritonRL shows only a slight increase (56% to 58%), suggesting it learned to generate genuine code.
Effectiveness of Hierarchical Reward Decomposition.
$r_{i}=\texttt{syntax}_{i}\cdot\texttt{func}_{i}\cdot(\beta\cdot\texttt{correct}_{i}+(1-\beta)\cdot\texttt{speedup}_{i}),$ (4)
where $\beta\in[0,1]$ is a hyperparameter that the ratio of correctness in the reward mixture. In Li et al. (2025b), the reward is computed only based on correctness conditioned and validity, which corresponds to $\beta=1.0$, while in Baronio et al. (2025), the reward is defined as the sum of correctness and speedup with some fixed weights. We defer the detailed GRPO formulation for uniform reward to Appendix E. We compare TritonRL using hierarchical reward decomposition against models trained with uniform reward designs for various $\beta$ values, keeping all other RL training settings (GRPO algorithm, hyperparameters, and pre-RL fine-tuning) consistent.
Reward Type Correctness ratio ($\beta$) valid compiled / correct $\textbf{fast}_{\textbf{1}}$ / $\textbf{fast}_{\textbf{2}}$
Hierarchical ($\alpha^{*}$) - $99.0$ $82.0$ / $\mathbf{56.0}$ $\mathbf{5.0}$ / $1.0$
Uniform $0.0$ $94.0$ $87.0$ / $47.0$ $3.0$ / $1.0$
Uniform $0.3$ $100.0$ $\mathbf{90.0}$ / $38.0$ $2.0$ / $1.0$
Uniform $0.5$ $92.0$ $77.0$ / $43.0$ $3.0$ / $1.0$
Uniform $0.7$ $98.0$ $78.0$ / $40.0$ $1.0$ / $1.0$
Uniform $1.0$ $95.0$ $75.0$ / $38.0$ $2.0$ / $1.0$
@@
Table 2: Comparison of hierarchical versus uniform reward designs on KernelBench Level 1 tasks. All metrics are reported as pass@10 (%). Reward type specifies whether the single reward is assigned uniformly across all tokens (Uniform), or the reward signal is decomposed by token class (Hierarchical) as in equation 3. The hyperparameter $\beta$ in the uniform reward design controls the proportion of correctness in the single reward function applied to all tokens. See equation 4 for details.
Effect of Plan-to-Code Update Ratio ($\alpha$).
We also analyze the effect of the hyperparameter $\alpha$ in our hierarchical reward decomposition, which controls the relative update rate of planning versus coding actions during RL training. Smaller $\alpha$ values slow down plan token updates, allowing coding actions to adapt to more stable planning distributions. As shown in Table 3, $\alpha=0.1$ achieves the best overall performance, while higher values (e.g., $\alpha=1.0$) lead to instability and reduced correctness and speedup. Conversely, setting $\alpha=0.0$, disabling plan updates, also degrades results, and this indicates that some plan adaptation is necessary. These findings highlight the importance of balancing plan and code updates to avoid premature convergence and maximize Triton code quality.
$0.0$ $98.0$ $78.0$ / $37.0$ $3.0$ / $1.0$
$0.1$ ($\alpha^{*}$) $99.0$ $82.0$ / $\mathbf{56.0}$ $\mathbf{5.0}$ / $1.0$
$0.3$ $100.0$ $\mathbf{96.0}$ / $55.0$ $1.0$ / $1.0$
$0.5$ $100.0$ $85.0$ / $48.0$ $2.0$ / $1.0$
$1.0$ $98.0$ $64.0$ / $33.0$ $2.0$ / $1.0$
Table 3: Ablation study of the hyperparameter $\alpha$ in TritonRL on KernelBench Level 1 tasks. All metrics are reported as pass@10 (%). $\alpha$ is the weighting factor for plan rewards in equation 3, determining how quickly plan tokens are updated relative to code tokens during training. The default configuration uses $\alpha^{*}=0.1$.
Effect of Data Mixture for RL Training.
We explore different data mixtures for RL training of TritonRL to understand how training data composition affects model performance. Our training dataset consists of two levels of tasks, with the ratio controlled by a data mixture probability vector $\bm{p}=[p_{1},p_{2}]$, where $p_{1}$ and $p_{2}$ represent the probabilities of sampling Level 1 and Level 2 tasks, respectively. We evaluate TritonRL on KernelBench Level 1 and 2 tasks using three data mixing strategies for RL training: training exclusively on Level 1 tasks ($\bm{p}=[1,0]$), exclusively on Level 2 tasks ($\bm{p}=[0,1]$), and a balanced mixture of both ($\bm{p}=[0.5,0.5]$), as summarized in Table 4. Training only on Level 1 tasks ($\bm{p}=[1,0]$) yields the best correctness and $\text{fast}_{1}$ on Level 1 evaluation, while training only on Level 2 tasks does not improve Level 2 performance. This may be because Level 2 tasks are inherently more complex, and reward signals from those tasks are very sparse, making it difficult for the model to learn effectively (as confirmed in Table 5). This observation motivates future work to explore adaptive $\bm{p}$ scheduling during post-training.
@@
Level 1 $[1,0]$ $99.0$ $82.0$ / $56.0$ $5.0$ / $1.0$ $66.0$ $29.0$ / $7.0$ $0.0$ / $0.0$
Level 1+2 $[0.5,0.5]$ $99.0$ $92.0$ / $43.0$ $2.0$ / $1.0$ $74.0$ $35.0$ / $8.0$ $0.0$ / $0.0$
Level 2 $[0,1]$ $100.0$ $97.0$ / $49.0$ $3.0$ / $1.0$ $57.0$ $37.0$ / $6.0$ $0.0$ / $0.0$
Table 4: Ablation study on data mixture for RL training of TritonRL, where the performance is evaluated on KernelBench level 1 and level 2 tasks.
Limited Performance on Fusion Tasks (Level 2).
We evaluated TritonRL and baseline models on KernelBench Level 2 tasks, which involve fused implementations such as Conv+Relu. As shown in Table 5, TritonRL outperforms other 8B-scale Triton-specific models in correctness and speedup. However, all models show a marked drop from Level 1 to Level 2, reflecting the increased difficulty of generating fully valid Triton code for fusion tasks. Fusion tasks require integrating multiple kernels, making it more challenging to produce correct implementations compared to single-operation replacements. Models often take shortcuts by implementing only part of the task in Triton and relying on high-level PyTorch APIs for the remainder, resulting in incomplete or invalid solutions (see Appendix F.3 for examples). Additionally, these tasks yield much sparser reward signals, since successful fusion implementations are rare, making effective learning more difficult even when similar fusion tasks are frequently encountered during training, as shown in Table 4. This pronounced performance gap highlights the need for improved optimization strategies and leaves substantial room for future progress.
Model #Params Level2 Level2 (w/o robust)
valid compiled / correct $\textbf{fast}_{1}$ / $\textbf{fast}_{2}$ compiled / correct
Qwen3 (base) 88B $56.0$ $1.0$ / $0.0$ $0.0$ / $0.0$ $52.0$ / $11.0$
Qwen3 1414B $35.0$ $24.0$ / $1.0$ $0.0$ / $0.0$ $94.0$ / $65.0$
Qwen3 3232B $31.0$ $16.0$ / $0.0$ $0.0$ / $0.0$ $73.0$ / $22.0$
KernelLLM 88B $0.0$ $0.0$ / $0.0$ $0.0$ / $0.0$ $96.0$ / $3.0$
AutoTriton 88B $70.0$ $3.0$ / $0.0$ $0.0$ / $0.0$ $97.0$ / $76.0$
TritonRL (ours) 88B $69.0$ $29.0$ / $\mathbf{7.0}$ $0.0$ / $0.0$ $88.0$ / $42.0$
   w/o RL (SFT only) 88B $67.0$ $\mathbf{32.0}$ / $6.0$ $0.0$ / $0.0$ $98.0$ / $41.0$
Claude-3.7 - $34.0$ $34.0$ / $\mathbf{12.0}$ $1.0$ / $0.0$ $98.0$ / $60.0$
Table 5: Main results on KernelBench Level 2. The left block reports evaluation with the robust verifier (syntax + functionality). The right block (w/o robust verifier) lacks functionality checks, leading to misleading correctness estimates, more severe than Level 1 evaluation.
@@ 4 Conclusion
In this work, we introduce TritonRL, a specialized LLM for Triton code generation, trained with a novel RL framework featuring robust verifiable rewards and hierarchical reward assignment. Our experiments on KernelBench show that TritonRL surpasses existing fine-tuned Triton models in validity, correctness, and efficiency. Ablation studies demonstrate that both robust reward design and hierarchical reward assignment are essential for achieving correctness and efficiency. We believe TritonRL marks a significant advancement toward fully automated and efficient GPU kernel generation with LLMs.
Reproducibility Statement
Our code-base is built upon publicly available frameworks (Verl (Sheng et al., 2025). Section 3.1 and the Appendix H G describe the experimental settings in detail.
Appendix A LLM Usage
We used an LLM to improve the writing by correcting grammar in our draft. It was not used to generate research ideas.
Appendix B Related Work
B.1 LLM for Kernel Generation
@@
Another line of research focuses on Triton kernel generation. Li et al. (2025a) introduced TritonBench, providing evaluations of LLMs on Triton programming tasks and highlighting the challenges of Triton’s domain-specific language and GPU programming complexity. To further enhance LLMs’ capabilities in Triton programming, Fisches et al. (2025) has introduced KernelLLM, a fine-tuned model of Llama3.1-8B-Instruct via supervised fine-tuning with Pytorch and Triton code pairs in KernelBook Paliskara & Saroufim (2025), but its performance is limited by the quality of training data. Similarly, Li et al. (2025b) introduced AutoTriton, a model fine-tuned specifically for Triton programming from Seed-Coder-8B-Reasoning Zhang et al. (2025), which achieves improved performance via SFT and RL with verifiable rewards based on correctness and rule-based Triton syntax verification, which may have limited improvement in runtime efficiency due to correctness-focused rewards. Both KernelLLM and AutoTriton are concurrent works developed alongside our work, and we provide a detailed comparison in Section 3.2.
B.2 Reinforcement Learning with Verifiable Rewards
@@ Appendix C Notations
  • $o_{i}$: $i$-th sample in the group $G$, which includes a reasoning trace that provides the ”plan” for Triton code optimization and implementation and the final ”Triton code”, i.e. $o_{i}=\{o_{i,\text{plan}},o_{i,\text{triton}}\}$
  • $T_{i}^{c}$: set of token indices corresponding to token class $c\in\{\text{plan},\text{triton}\}$ in the $i$-th sample
  • $r^{c}(q,o_{i})=r_{i}^{c}$: reward function for token class $c\in\{\text{plan},\text{triton}\}$.
  • $\hat{A}_{t}^{c}(q,o_{i})=r^{c}(q,o_{i})-\frac{1}{G}\sum_{j=1}^{G}r^{c}(q,o_{j})$: token-level advantage of the $t$-th token of the $i$-th sample belonging to token class $c\in\{\text{plan},\text{triton}\}$, shortened as $\hat{A}_{i,t}^{c}$.
Appendix D Metrics
We provide the formal definitions of the evaluation metrics used in this paper. Given a set of $N$ tasks $\{q_{n}\}_{n=1}^{N}$ and $k$ samples $\{o_{i}\}_{i=1}^{k}$ generated by the model for each task, we define the following metrics:

### v2

@@ 3.2 Main Experiment Results
Correctness and Speedup Evaluation on KernelBench.
[…]
In Appendix F.1, we present additional evaluation results across varying $k$ values for pass@$k$ metrics. These results demonstrate that TritonRL consistently outperforms baselines as $k$ increases, indicating that our model generates more diverse candidate solutions and effectively leverages additional sampling during inference.
@@
Table 2: Comparison of hierarchical versus uniform reward designs on KernelBench Level 1 tasks. All metrics are reported as pass@10 (%). Reward type specifies whether the single reward is assigned uniformly across all tokens (Uniform) as in (3), or the reward signal is decomposed by token class (Hierarchical) as in (2). The hyperparameter $\beta$ controls the proportion of correctness in the single reward function applied to all tokens. The models are trained from the same SFT checkpoint distilled from DeepSeek-R1.
Reward Type $\bm{\beta}$ valid compiled / correct $\textbf{fast}_{\textbf{1}}$ / $\textbf{fast}_{\textbf{2}}$
Hierarchical ($\alpha^{*}$) - $100.0$ $100.0$ / $\mathbf{78.0}$ $\mathbf{36.0}$ / $13.0$
Uniform $0.0$ $100.0$ $100.0$ / $65.0$ $33.0$ / $\mathbf{15.0}$
Uniform $0.3$ $100.0$ $100.0$ / $68.0$ $29.0$ / $14.0$
Uniform $0.5$ $100.0$ $100.0$ / $68.0$ $35.0$ / $13.0$
Uniform $0.7$ $100.0$ $100.0$ / $70.0$ $34.0$ / $12.0$
Uniform $0.9$ $100.0$ $100.0$ / $75.0$ $31.0$ / $11.0$
Uniform $1.0$ $100.0$ $100.0$ / $70.0$ $35.0$ / $12.0$
Table 3: Ablation study on data mixture for RL training of TritonRL, where the performance is evaluated on KernelBench level 1 and level 2 tasks. The models are trained from the same SFT checkpoint distilled from DeepSeek-R1.
@@
Level 1 $[1,0]$ $100.0$ $\mathbf{100.0}$ / $\mathbf{78.0}$ $\mathbf{36.0}$ / $13.0$ $\mathbf{63.0}$ $\mathbf{63.0}$ / $\mathbf{23.0}$ $\mathbf{12.0}$ / $3.0$
Level 1+2 $[0.5,0.5]$ $100.0$ $99.0$ / $73.0$ $33.0$ / $\mathbf{14.0}$ $\mathbf{63.0}$ $\mathbf{63.0}$ / $22.0$ $9.0$ / $3.0$
Level 2 $[0,1]$ $100.0$ $\mathbf{100.0}$ / $70.0$ $30.0$ / $10.0$ $50.0$ $50.0$ / $16.0$ $9.0$ / $\mathbf{4.0}$
Robust Verification Prevents Reward Hacking.
In Figure 3, we analyze the validity of Triton codes generated by fine-tuned models to understand the types of errors each model is prone to. Although TritonRL and other fine-tuned baselines (KernelLLM and AutoTriton) learn to generate more valid codes after fine-tuning, a more detailed breakdown reveals that they exhibit a much higher proportion of functionally invalid codes (red). In contrast, TritonRL generates significantly fewer invalid codes in terms of both syntax and functionality after fine-tuning, demonstrating the effectiveness of our robust verification in enhancing code quality without reward hacking.
Moreover, Table 1 highlights how heavily prior models relied on cheating shortcuts. Without functionality verification (i.e., w/o the robust verifier), AutoTriton’s correctness jumps from 57% to 87%, revealing its tendency to exploit loopholes by superficially meeting syntax checks rather than producing truly functional Triton codes. In contrast, TritonRL shows only a slight increase ($\leq 3$%), suggesting it learns to generate genuine code without relying on shortcuts.
Effectiveness of HRD: Hierarchical vs. Uniform.
$\begin{aligned} r_{i,j}&=\beta\cdot R^{\text{correct}}(g_{i},o_{i,j})+(1-\beta)\cdot R^{\text{speedup}}(g_{i},o_{i,j}),\end{aligned}$
where $\beta\in[0,1]$ is a hyperparameter that controls the ratio of correctness in the reward mixture. In Li et al. (2025b), the reward is computed only based on correctness, which corresponds to $\beta=1.0$, while in Baronio et al. (2025), the reward is defined as the sum of correctness and speedup with some fixed weights. We defer the detailed GRPO formulation for uniform reward to Appendix D. We compare TritonRL using hierarchical reward decomposition against models trained with uniform reward designs for various $\beta$ values, keeping all other RL training settings (GRPO algorithm, hyperparameters, and pre-RL fine-tuning) consistent.
Ablation Study on Plan-to-Code Update Ratio ($\alpha$).
We also analyze the effect of the hyperparameter $\alpha$, which controls the relative update rate of planning versus coding actions during RL training. Smaller $\alpha$ values slow down plan token updates, allowing coding actions to adapt to more stable planning distributions. As shown in Table 4, $\alpha=0.1$ achieves the best overall performance, while higher values (e.g., $\alpha=1.0$) lead to instability and reduce correctness and speedup. Conversely, setting $\alpha=0.0$, disabling plan updates, also degrades results, indicating that some plan adaptation is necessary. These findings highlight the importance of balancing plan and code updates to avoid premature convergence and maximize Triton code quality.
Table 4: Ablation study of the hyperparameter $\alpha$ in TritonRL on KernelBench Level 1 tasks. All metrics are reported as pass@10 (%). $\alpha$ is the weighting factor, determining how quickly plan tokens are updated relative to code tokens during training. The default configuration uses $\alpha^{*}=0.1$ and the models are trained from the same SFT checkpoint distilled from DeepSeek-R1.
$0.0$ $100.0$ $100.0$ /$72.0$ $32.0$ / $12.0$
$0.1$ ($\alpha^{*}$) $100.0$ $100.0$ / $\mathbf{78.0}$ $\mathbf{36.0}$ / $13.0$
$0.5$ $100.0$ $100.0$ / $75.0$ $35.0$ / $12.0$
$0.8$ $100.0$ $99.0$ / $70.0$ $33.0$ / $10.0$
$1.0$ $100.0$ $100.0$ / $71.0$ $31.0$ / $11.0$
Comparison of Difficulty-Based Data Mixtures.
We explore different data mixtures for RL training of TritonRL to understand how training data composition affects model performance. Our training dataset consists of two levels of tasks, with the ratio controlled by a data mixture probability vector $\bm{p}=[p_{1},p_{2}]$, where $p_{1}$ and $p_{2}$ represent the probabilities of sampling Level 1 and Level 2 tasks, respectively. We evaluate TritonRL on KernelBench Level 1 and 2 tasks using three data mixing strategies for RL training: training exclusively on Level 1 tasks ($\bm{p}=[1,0]$), exclusively on Level 2 tasks ($\bm{p}=[0,1]$), and a balanced mixture of both ($\bm{p}=[0.5,0.5]$), as summarized in Table 3. Interestingly, training exclusively on Level 1 tasks ($\bm{p}=[1,0]$) yields the best correctness and $\text{fast}_{1}$ on both Level 1 and Level 2 evaluations. Including Level 2 tasks during RL training provides only marginal improvements or worsens performance on Level 2 tasks. This may be because Level 2 tasks are inherently more complex, and reward signals from those tasks are very sparse, making it difficult for the model to learn effectively. Consequently, it suggests that focusing on fundamental single-operation tasks builds more robust Triton block generation skills that transfer effectively to both simple and complex fusion scenarios, rather than directly training on the more challenging fusion problems themselves.
Diverse Input Augmentation Improves Triton Quality.
To assess the impact of input augmentation (IA) used for RL data construction (Section 2.1.2), we compare TritonRL trained with and without IA under identical settings in Table 1. Especially, TritonRL (G) + IA substantially improves over TritonRL (G) without IA on both Level 1 and Level 2 tasks, correctness increases by $12\%$ and fast1 by $6\%$ on KernelBench Level 1. This demonstrates that IA diversifies the training data and helps the model generalize to produce higher-quality Triton kernel codes. To demonstrate robustness to diverse input shapes, we evaluate TritonRL with IA on augmented KernelBench tasks with varied input shapes (see Appendix F.2).
Challenges and Future Directions.
Table 1 demonstrates that TritonRL maintains its advantage over both much larger (>>100B) models and 8B-scale Triton-specific models on KernelBench Level 2 fusion tasks. However, consistent with prior work (Li et al., 2025b; Fisches et al., 2025), fusion task generation remains a significant open challenge for all models, with relatively lower performance compared to Level 1 single-kernel tasks. Our analysis reveals two primary factors: (1) models frequently produce partial implementations that delegate computation to high-level PyTorch APIs, and (2) the sparsity of successful fusion implementations leads to weak reward signals during training (Table 3). Despite these challenges, TritonRL achieves the best results among models of comparable or larger scale, and our RL framework, with its robust verification and hierarchical reward mechanisms, provides essential foundations for future work in complex kernel generation. As kernel patterns grow more diverse, extending the verifier with formal guarantees, evaluating across additional benchmarks and accelerator targets, and exploring how HRD interacts with larger model scales represent natural next steps toward a more general hardware-aware RL paradigm.
@@ 4 Conclusion
In this work, we introduce TritonRL, a specialized LLM for Triton code generation, trained with a novel RL framework featuring robust verifiable rewards and hierarchical reward assignment. Our experiments on KernelBench show that TritonRL surpasses existing fine-tuned Triton models in validity, correctness, and efficiency. Ablation studies demonstrate that both robust reward design and hierarchical reward assignment are essential for achieving correctness and efficiency.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Related Work
A.1 LLM for Kernel Generation
@@
Another line of research focuses on Triton kernel generation. Li et al. (2025a) introduced TritonBench, providing evaluations of LLMs on Triton programming tasks and highlighting the challenges of Triton’s domain-specific language and GPU programming complexity. To further enhance LLMs’ capabilities in Triton programming, Fisches et al. (2025) have introduced KernelLLM, a fine-tuned model of Llama3.1-8B-Instruct via supervised fine-tuning with Pytorch and Triton code pairs in KernelBook Paliskara and Saroufim (2025), but its performance is limited by the quality of training data. Similarly, Li et al. (2025b) introduced AutoTriton, a model fine-tuned specifically for Triton programming from Seed-Coder-8B-Reasoning Zhang et al. (2025), which achieves improved performance via SFT and RL with verifiable rewards based on correctness and rule-based Triton syntax verification, which may have limited improvement in runtime efficiency due to correctness-focused rewards. Both KernelLLM and AutoTriton are concurrent works developed alongside our work, and we provide a detailed comparison in Section 3.2.
A.2 Reinforcement Learning with Verifiable Rewards
@@ Appendix B Notations
  • $o_{i,j}$: $j$-th sample in the group $G$ for the $i$-th prompt, which includes a reasoning trace that provides the ”plan” for Triton code optimization and implementation and the final ”Triton code”, i.e. $o_{i,j}=\{o_{i,j}^{\text{plan}},o_{i,j}^{\text{code}}\}$
  • $T_{i,j}^{c}$: set of token indices corresponding to token class $c\in\{\text{plan},\text{code}\}$ in the $j$-th sample of the $i$-th prompt
  • $r_{i,j}^{c}$: reward function for token class $c\in\{\text{plan},\text{code}\}$.
  • $A_{i,j}^{c}=r_{i,j}^{c}-\frac{1}{G}\sum_{j=1}^{G}r_{i,j}^{c}$: token-level advantage of the $j$-th sample for prompt $q_{i}$ belonging to token class $c\in\{\text{plan},\text{code}\}$.
Appendix C Metrics
We provide the formal definitions of the evaluation metrics used in this paper. Given a set of $N$ tasks $\{g_{n}\}_{n=1}^{N}$ and $k$ samples $\{o_{n,i}\}_{i=1}^{k}$ generated by the model for each task, we define the following metrics: